OpenStudy (anonymous):

Tyrone has a total of 50 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have?

OpenStudy (anonymous):

38

OpenStudy (anonymous):

x + y = 50 x = 3y + 2 3y + 2 + y = 50 4y + 2 = 50 4y = 48 y = 12 x = 38 CD = 38 DVD = 12

OpenStudy (anonymous):

sorry. I read your question wrong the first time.

OpenStudy (anonymous):

Its ok i knew it couldnt be 152 i just figured you read it wrong

OpenStudy (anonymous):

no of dvd's=12 no of cd's=38